Use Laplace transform to identify F(x)

y''-y'-6y=0 with y(0)=1 & y'(0)=2
Answer #1

Answer 0 y-yl-6y=0 =0 , yco)=1 and y(o) = 2 We have L[y(t)} = 5² 4,5) - syco)-ylo L[y(4)] = sy.s)-90) Leat - no (3-a)nt Y=> A+B=1 @ b -3A + 2B=1 2x@-0 24+2B = 2 -3A +2B =) 5A A = 75 .415 Theis B=1-À =1-% = Then ② => 415 + 5+2 5-3 Thers, use the i